OpenStudy (anonymous):

A child and a sled with a combined mass of 50.0kg slide down a frictionless slope. If the sled starts from rest and has a speed of 3.00m/s at the bottom, what is the height of the hill?

OpenStudy (gollywolly):

change in kinetic energy is going to equal change in gravitational potential energy.\[\Delta K=1/2 mv^2\] \[\Delta U_g=mg \Delta x\]

OpenStudy (anonymous):

Which numbers would I plug into the equation @gollywolly

OpenStudy (gollywolly):

\[1/2(50.0)(3.00)^2=(50.0)(9.81)\Delta x\]

OpenStudy (anonymous):

Thank you so much! @gollywolly

OpenStudy (gollywolly):

No problem. Do you understand why that works? @daisy_arlene

OpenStudy (anonymous):

Yeah, somewhat @gollywolly

OpenStudy (shamim):

U can solve this problem in another way

OpenStudy (shamim):

v^2=u^2+2gh

OpenStudy (shamim):

h=?

OpenStudy (shamim):

Given u=0

OpenStudy (shamim):

v=3m/s